I would like to aknwoledge the participation, contribution/support from:

• Gordon Bromage and Hannah Leigh Worters, UCLAN, UK

• Carl Foley, Hilary McKay, Urmila Mitra Kraev, MSSL, UK

• Alan Whiting group (CTIO, Chile), in particular Melissa Rice.

• Carlos Alberto, Germano Quast and Ramiro Delareza (Brasil).

• The XMM/SOC at Villafranca, Madrid (Spain), for their professional support. In particular, of Nora Loiseau and Pedro M. Rodriguez Pascual.

• Federico Gonzalez (CASLEO), Argentina

• Helen Mason, University of Cambridge UK

2RE J0241-53 (nicknamed Horace) is a binary of active X-ray emitting stars, classified as a member of an association of young Post-T-Tauri stars (∼ 30 Myr) by Torres et al. (2000).

Goal: determine the physical characteristics of the X-ray emission (during quies-

cent and flaring state) of the visual binary 2RE J0241-53, and relate them to the properties of the stellar photospheres.

In particular in terms of chemical composition and understand the evolution from the T-Tauri phase.

It is likely that the X-ray emission from CTTS is due to accretion (as proposed by Kastner et al. 2002 for TW Hya), while the PTTS have extended solar-like coronae.

The target 2RE J0241-53 (EUVE J0241-530) is a bright XUV visual binary, discovered

by the ROSAT (Pye et al., 1995) and EUVE (Bowyer et al., 1996) all-sky surveys (see also Jeffries et al. 1996). Count rates in the 100 Å bandpassess were comparable to those of the brightest flare stars.

Both stars have strong chromospheric emission and have recurrent flares (Ball and Bromage 1995).

dK7e and dM3e young and fast rotating stars -Southern star is Li-rich. Pushing to the limits

- OM fast mode windows -angular spearation of 22”. -multi-continent, multi-cultural experience Ground-based spectroscopy

• SAAO 1.9m grating spectrograph with 1 Å resolution and ≃90 s exposures.

• Wavelength range: 3700 - 4360 Å • 6 consecutive nights Ground-based photometry

• UBV aperture photometry with the SAAO 0.5-m telescope

• 3 nights On XMM night

• CASLEO: cloudy

• SAAO 1.9m spectra 18:32 UT – 00:04 UT

• CTIO low-resolution (17.2 Å FWHM) 1.5m spectra

• CTIO CCD B and U band photometry with the 0.9m 01:17 04:40 UT • SAAO U-band monitoring

• Brasil: B-band monitoring 23:00 – 00:30 UT

- two ground station handovers, one on 10-02-2004 at 04:14, and 05:37 UT. During handovers, that have a margin of about 1800 sec., the OM has to remain closed, which creates gaps in the OM data. The handovers do not affect the other instruments.

1 Summary of preliminary results

• This is one of the few (two ?) X-ray observations of PTT stars.

• The XMM/OM fast-mode windows worked very well but data are missing.

• Both stars confirmed to be active, with flares.

• Ground-based SAAO and CTIO observations are nicely complementary.

• We confirm Linsky’s law (the most interesting events occurr at the end of the observation).

Figure 1: SAAO spectra of the northern star - 7th and 9th (XMM night) Feb 2004 and of the southern star - 7th Feb 2004.

4 Figure 2: Bottom two plots: consecutive 90s exposures from XMM night (19:12,19:14 UT). Notice the variation in the Ca II (H and K) and H δ.

6 Figure 3: One XMM/OM optical image (full FOV, UVW1 filter) of the target. Below are two sample images from the fast-mode extraction windows.

Figure 4: XMM OM photometry 8 Figure 5: One EPIC MOS 1 image

9 Figure 6: One EPIC PN light curves 10 Figure 7: Flares on Northern star.

11 Figure 8: Flares on the Southern star.

12 Figure 9: EPIC MOS 1 spectrum of the Southern star.

13 Figure 10: Top: the RGS 1 first order predicted spectrum. Bottom: the observed one. Where is all the Iron gone ?

Figure 11: Horace vs. PZ Tel 15 Figure 12: Ne IX spectral region.

16 Figure 13: O VII spectral region.

17